OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 10 inches. What is the volume of water in the vase?

OpenStudy (anonymous):

volume of water=vol of cylinder till 10 inc-vol of the 6 marbles

OpenStudy (anonymous):

But I'm confused on how to find the volume of the marbles?

OpenStudy (mertsj):

The marbles are spheres. The volume of a sphere is \[V=\frac{4}{3}\pi r^3\]

OpenStudy (mertsj):

There are 6 marbles and their radius is 1 so replace r with 1 in the formula and multiply the answer by 6 since there are 6 marbles. That will be the volume of the marbles.

OpenStudy (anonymous):

so the volume of the marbles would end up becoming 75.36

OpenStudy (mertsj):

\[V _{m}=6(\frac{4}{3})\pi (1)^3=8\pi \]

OpenStudy (anonymous):

Would i divide the 42.39 that i got by the 25.12 i got with the 8\[\pi\] ?

OpenStudy (mertsj):